# A new foundation of quantum decision theory

### v4

@@ 3 Quantum decision theory in practice; an example
$P(A\ \mathrm{helps})$, $P(B\ \mathrm{helps})$,
$P(A\ \mathrm{helps\ and}\ B\mathrm{\ helps\ given\ that\ he\ knows\ that\ }A\mathrm{\ helps})$.
@@
$P(A\ \mathrm{helps\ and\ also}\ B\mathrm{\ helps})=P(B\ \mathrm{helps\ and\ also\ }A\mathrm{\ helps})$
$=P(A\ \mathrm{helps})P(B\ \mathrm{helps}).$
@@
$P(A\ \mathrm{helps\ and\ also}\ B\mathrm{\ helps})=$
$P(A\ \mathrm{helps})P(B\ \mathrm{helps}|A\ \mathrm{helps})=$
@@ 3.1 Some consequences
The marginal probability that $B$ helps is $P(B\ \mathrm{helps})=\|\Pi_{B}|\psi\rangle\|^{2}$. Then by a geometric argument it can be seen that there exist situations where
$P(A\ \mathrm{helps\ and\ also}\ B\mathrm{\ helps})>P(B\ \mathrm{helps}),\mathrm{\ i.e.,}$
@@
$P(A\ \mathrm{helps\ and\ also}\ B\mathrm{\ helps})+P(A\ \mathrm{helps\ but}\ B\mathrm{\ does\ not\ help})$
$=|a|^{2}+|b|^{2}.$
@@
$P(A\ \mathrm{helps\ and\ also}\ B\mathrm{\ helps})\neq P(B\ \mathrm{helps\ and\ also\ }A\mathrm{\ helps}).$
@@ 4.1 A summary of assumptions
It is crucial that this article relies upon my earlier work on the foundation of quantum theory. This work in turn relies on several assumptions, as stated here in connection to the discussion around Theorem 1 in Subsection 2.1:
1) The variables, here decision variables are divided into accessible and inaccessible ones. A decision variable is said to be accessible to a person $C$ if and only if $C$ is able to carry out the corresponding decision.
2) All accessible decision variables can be seen as functions of some large inaccessible theoretical variable. In practice this may be taken to mean that all earlier experiences of $C$, among other things related to his earlier decisions, are collected as a part of an hypothetical large variable $\phi$. Note that $C$ has not grown up in isolation. He is the result of interaction with other people, in particular the result of a specific human culture. Depending upon circumstances, he may have made specific conscious decisions to free himself from this culture, but nevertheless, the background is there.
@@
In addition to these 4 assumptions, this article also relies on specific assumptions behind the Born rule, the means for calculating quantum probabilities -see Subsections 2.4-2.6:
@@
What are then my own ideals? During my work with the quantum foundation, I have had a strong religious faith, as discussed explicitly in the two articles Helland (2022d) and Helland (2023d). I do believe in the existence of a God, and I have sympathies with the best sides of all religious beliefs. Personally, I have chosen to be a Christain, but I do well understand people that have made other choices here. Religion is definitely connected to culture.
@@ 4.3 Everyday decisions
The exception to this use of Born’s rule in a decision context, may be what Kahneman (2011) calls decisions based upon slow thinking, decision where the actor $C$ actively uses his conscious self in the process of making the decisions. It is very unlikely that such decisions can be modeled by QDT.
@@ 4.4 Decisions made by political leaders
Many of the issues that political leaders have to make decision on, are really difficult, often too difficult to be left to a single person. In democratic countries, joint decisions are often made by a group of communication persons. Such joint discussions are not discussed in the present article, but they could have been, in the light of a version of Theorem 1 which is valid for a group of communicating actors, see Helland (2024a).
@@
Certain political leaders may make decisions that really have cruel consequences. Some examples are 1) the cruel attack on Ukraine initiated by Vlademir Putin in February 2022; 2) the attack on innocent sivilians ordered by the leaders of Hamas in October 2023; 3) the continuous attacks on innocent people in Gaza initiated by Israel’s leaders.
@@
One clear wish is that they never should be used again. How can that be avoided? My idea here is that this may be connected to the observation that all people, also political leaders may have some religious feelings, at least some thoughts around their destiny after death. May be the following thought could be spread and made common: The political leader who ever dares to order some atomic bomb to be released, deserves an extremely hard punishment after death. If this could be believed by all our leaders, the atomic arsenal spread around the world will be totally useless, and negotiations around the reduction of this arsenal might start. This is of course an idealistic thought, but on this important issue, one should really try to think in a new way.
@@ 5 Concluding remarks
In this article I have concentrated on decisions made by a single person. But the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two variables are here called complementary if they are really different (not in one-to-one correspondence) and maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
@@
It is a hope that ideas related to this article may lead to an increased focus on making the best decisions for all people, for political leaders and in a global perspective, the best decisions in order to obtain peace on this earth and a viable future for all human beings.
@@
I am grateful to Solve Sæbø and Trygve Almøy for discussions.

### v5

@@ 3 Quantum decision theory in practice; an example
P(Ahelps)P(A\ $\mathrm${helps}), P(Bhelps)P(B\ $\mathrm${helps}),
[…]
@@
[…]
=P(Ahelps)P(Bhelps).=P(A\ $\mathrm${helps})P(B\ $\mathrm${helps}).
@@
[…]
P(Ahelps)P(Bhelps|Ahelps)=P(A\ $\mathrm${helps})P(B\ $\mathrm${helps}|A\ $\mathrm${helps})=
@@ 3.1 Some consequences
[…]
@@
[…]
$|a|^{2}+|b|^{2}.$
@@
[…]
@@ 4.1 A summary of assumptions
It is crucial that this article relies upon my earlier work on the foundation of quantum theory. This work in turn relies on several assumptions, as stated here in connection to the discussion around Theorem 1 in Chapter 2:
1) The variables, here decision variables, are divided into accessible and inaccessible ones. A decision variable is said to be accessible to a person $C$ if and only if $C$ is able to carry out the corresponding decision.
2) All accessible decision variables can be seen as functions of some large inaccessible theoretical variable. In practice this may be taken to mean that all earlier experiences of $C$, among other things related to his earlier decisions, together with what may be called his free will, are collected as a part of an hypothetical large variable $\phi$. Note that $C$ has not grown up in isolation. He is the result of interaction with other people, in particular the result of a specific human culture. Depending upon circumstances, he may have made specific conscious decisions to free himself from this culture, but nevertheless, the background is there.
@@
In addition to these 4 assumptions, this article also relies on specific assumptions behind the Born rule, the means for calculating quantum probabilities -see Chapter 2 and Chapter 3:
@@
What are then my own ideals? During my work with the quantum foundation, I have had a strong religious faith, as discussed explicitly earlier in this book and in the two articles Helland (2022d) and Helland (2023d). I do believe in the existence of a God, and I have sympathies with the best sides of all religious beliefs. Personally, I have chosen to be a Christain, but I do well understand people that have made other choices here. Religion is definitely connected to culture.
@@ 4.3 Everyday decisions
The exception to this use of Born’s rule in a decision context, may be what Kahneman (2011) calls decisions based upon slow thinking, decision where the actor $C$ actively uses his conscious self in the process of making the decisions. It seems unlikely that such decisions can be simply modeled by QDT.
@@ 4.4 Decisions made by political leaders
Many of the issues that political leaders have to make decision on, are really difficult, often too difficult to be left to a single person. In democratic countries, joint decisions are often made by a group of communication persons. Such joint discussions are not discussed in the present chapter, but they could have been, in the light of a version of Theorem 1 which is valid for a group of communicating actors, see Helland (2024a).
@@
Certain political leaders have made decisions that really have cruel consequences. Some examples are
1) The cruel attack on Ukraine initiated by Vlademir Putin in February 2022. The war is continuing here, and through propaganda and other means, many Russians support the war. Putin and his followers seem to base much of their motivation on the notion of power and on mistaken judgements. Vladimar Putin claims to be a Christian, but he seems to have forgotten Christ’s main command, urging us to love all humans in addition to loving God.
2) The attack on innocent sivilians ordered by the leaders of Hamas in October 2023;
3) The continuous attacks on innocent people in Gaza initiated by Israel’s leaders. This has escalated now to a terrible humanitarian situation. Benjamin Netanyahu and his rightwing gouvernment are the main responsible here, but unfortunately, in some of their decisions they are supported by a substancial part of Israel’s population.
4) The decisions taken by generals on the both sides in the confict in Sudan. These decisions have resulted in a quite unnecessary civilian war, which has implied terrible suffering for a large part of Sudan’s population.
It is a hope that these leaders in due time will repent and realize that we all, in our final phase, may be subject to a divine judgement.
@@
One clear wish is that they never should be used again. How can that be avoided? My idea here is that this may be connected to the observation that all people, also political leaders may have some religious feelings, at least some thoughts around their destiny after death. May be the following thought could be spread and made common: The political leader who ever dares to order some atomic bomb to be released over humans, deserves an extremely hard punishment after death. If this could be believed by all our leaders, the atomic arsenal spread around the world will be totally useless, and negotiations around the reduction of this arsenal might start. This is of course an idealistic thought, but on this important issue, one should really try to think in a new way.
@@ 5 Concluding remarks
In this article, I have concentrated on decisions made by a single person. But the theory can be extended to joint decisions made by a group of communicating persons. Then these persons must have common ideals, as modelled by a perfectly rational higher being $D$. The ideals must be common in the setting determined by the relevant decision variables.
@@
As a basis for quantum decision theory as discussed here, lies the idea of two complementary decision variables at the same time in the mind of some person or in the joint minds of a group of communicating persons. I repeat that two variables are here called complementary if they are really different (not in one-to-one correspondence), have similar ranges, and are maximal as accessible variables. Another essential assumption is the existence of ideals that are perceived as perfectly rational, modelled by a hypothetical ‘ideal higher being’.
@@
It is a hope that ideas related to this chapter may lead to an increased focus on making the best decisions for all people, for those given by political leaders in a global perspective, the best decisions in order to obtain peace on this earth and a viable future for all human beings.
@@
I first want to thank Solve Sæbø and Trygve Almøy for discussions. This article is partly based upon Chapter 11 in Helland (2025c). I am grateful for permission given by the editors here.